OpenStudy (anonymous):

Given a = 4 and b = -2, evaluate -|b - a|. i got -6?

OpenStudy (anonymous):

ryt

OpenStudy (anonymous):

I am not sure but I think yes.

OpenStudy (lalaly):

|b-a| =b-a and -(b-a) so -|b-a|= -(b-a) and b-a so u get 6 and -6